# Dynamics of non-Hermitian Floquet Wannier-Stark system

### v1

@@ Abstract
We study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation, where the hopping strength is a periodic function of time with Floquet frequency $\mathcal{\omega}$. It is shown that the energy level of the instantaneous Hamiltonian is still equally spaced and independent of time $t$ and the Hermiticity of the hopping term. In the case of off resonance, the dynamics are still periodic, while the occupied energy levels spread out at the resonance, exhibiting $t^{z}$ behavior. Analytic analysis and numerical simulation show that the level-spreading dynamics for real and complex hopping strengths exhibit distinct behaviors and are well described by the dynamical exponents $z=1$ and $z=1/2$, respectively.
I Introduction
The Hermitian and static system is the main study object of traditional quantum mechanics, preserving the reality and conservation of energy. When these two constraints are violated, the properties of the material could be drastically modified, leading to new phases and phenomena Moiseyev2011 ; Tannor2007 without analogs in their static or Hermitian counterparts. Floquet engineering has facilitated the exploration of diverse dynamic, topological, and transport phenomena across a wide range of physical contexts. More recently, the Floquet approach has been employed to engineer topological phases in non-Hermitian systems. Importantly, it was discovered that the interplay between time-periodic driving fields and the presence of gain, loss, or nonreciprocal effects can lead to the emergence of topological phases exclusive to non-Hermitian Floquet systems PhysRevB.98.205417 ; PhysRevLett.102.065703 ; LeonMontiel2018 ; Li2019 ; PhysRevA.102.062201 ; Blose2020 ; Zhou2020 ; Zhang2020 ; Xiao2020 ; Ageev2021 . Motivated by the above investigations, the aim of the present paper is to examine how non-Hermiticity and periodic driving impact a Stark ladder system. Specifically, we study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation. In this work, we introduce non-Hermiticity and periodic driving by focusing on cosinusoidal time-dependent hopping strength with frequency $\mathcal{\ \omega}$ and complex amplitude. The main reason for choosing such systems is that the energy level of the instantaneous Hamiltonian is still equally spaced with $\mathcal{\omega}_{0}$ and independent of time $t$ and the Hermiticity of the hopping strength. It allows us to investigate the dynamics based on analytical solutions. We find that, in the case of off resonance, the dynamics are still periodic, while the occupied energy levels spread out at the resonance, exhibiting $t^{z}$ behavior. Analytic analysis and numerical simulation show that the level-spreading dynamics for real and complex hopping strengths exhibit distinct behaviors and are well described by the dynamical exponents $z=1$ and $z=1/2$, corresponding to superdiffusion and diffusion, respectivelyZhao2006 ; metzler1999anomalous ; Bouchaud1990 . These findings may help us deepen our understanding of the interplay between non-Hermiticity and periodic driving impacts on the dynamics of the system. In addition, we propose a scheme to demonstrate the results by wavepacket dynamics in a $2$D square lattice. This paper is organized as follows. In Sec. II, we introduce a non-Hermitian time-dependent tight-binding model supporting a Wannier-Stark ladder and present the analytical eigenstates of the instantaneous Hamiltonian. In Sec. III, we focus on the dynamics of the static system, while the Floquet system at resonance in Sec. IV. Sec. V describes the application of the solutions to a specific initial state. In Sec. VI, a scheme for the simulator of our results is proposed in the framework of wavepacket dynamics in a $2$D square lattice. Finally, we summarize our results in Sec. VII.
@@ I Introduction
Figure 1: (a) Schematic illustrations of the Hamiltonian in Eq. (1), which represents a tight-binding chain with a time dependence on the hopping strength between two adjacent sites. The on-site potential is tilted with slope $\omega_{0}$. It is shown that such a system supports a Stark-ladder with time-independent energy-level spacing. (b) Schematic illustrations for the Hamiltonian in Eq. (43), which is a tight-binding model of a square lattice with time-independent system parameters. The contributions of hopping strength and on-site potential can be used to simulate the dynamics of system (a).
II Model and generalized Wannier-Stark ladder
@@
where $|n\rangle$ denotes a site state describing the Wannier state localized on the $n$th period of the potential. Here, $\kappa(t)$ is the time-dependent tunneling strength, while $\mathcal{\omega}_{0}$ is the slope of the linear potential from a static force. Fig. 1(a) shows a schematic of the system. In our study, $\kappa$ can be real and complex. It is well known that the spectrum of $H$ is equally spaced when $\kappa$ is a real constant, supporting periodic dynamics Hartmann2004 ; bloch1928elektron ; 1934 . This has been observed in systems of semiconductor superlattices and ultracold atoms Waschke1993 ; BenDahan1996 ; Wilkinson1996 ; Anderson1998 . Now, we extend the conclusion to the case with arbitrary $\kappa$. The eigenstate of $\left|\psi_{n}\right\rangle$ can always be written in the form
$\left|\psi_{n}\right\rangle=\sum_{m}c_{mn}|m\rangle,$ (2)
satisfying the Schrodinger equation
$H\left|\psi_{n}\right\rangle=E_{n}\left|\psi_{n}\right\rangle.$ (3)
The coefficient $c_{mn}$ is determined by the equation
$c_{m+1,n}+c_{m-1,n}=\frac{E_{n}-m\mathcal{\omega}_{0}}{\kappa}c_{mn},$ (4)
which accords with the recurrence relation of the Bessel function $J_{m}$ of argument $x$, i.e.,
$J_{m+1}(x)+J_{m-1}(x)=2\frac{m}{x}J_{m}(x).$ (5)
@@
$c_{mn}=J_{m-n}(-\frac{2\kappa}{\omega_{0}}),E_{n}=n\omega_{0}.$ (6)
We can see that the eigenenergy $n\mathcal{\omega}_{0}$ is independent of $\kappa$, regardless of whether it is real or complex. This is crucial for the main conclusion of this work. Although the profile of the eigenstate
@@
$\text{IPR}_{m}=\frac{\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{4}}{\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{2}}=\sum_{n}|J_{n-m}(-\frac{2\kappa}{\omega_{0}})|^{4}.$ (8)
@@
Figure 2: Plots of $P_{n}(t)$ and $P(t)$ defined in Eqs. (15 ) and (16), obtained by numerical diagonalization for several typical values of system parameters with $\mathcal{\omega}_{0}=1$. Here, (a1), (b1), and (c1) correspond to the cases with $\omega=0$, $0.01$, and $0.1$, respectively, and $\kappa_{0}=1$. (a2), (b2), and (c2) correspond to the cases with $\omega=0$, $0.01$, and $0.1$, respectively, and $\kappa_{0}=i$. (a3), (b3),and (c3) correspond to the cases with $\omega=0$, $0.01$, and $0.1$, respectively, and $\kappa_{0}=e^{i\pi/4}$.
III Bloch oscillation in non-Hermitian system
Now, we consider the dynamics of the system with slowly varying $\kappa(t)$. When $\kappa$ is time independent, the dynamics of the system are governed by the propagator
@@ III Bloch oscillation in non-Hermitian system
the Dirac probability distribution in real space for the evolved state $\left|\Phi(t)\right\rangle$ is
@@
which is still periodic. We obtain the conclusion that the total Dirac probability
@@
is still periodic, regardless of whether $\kappa$ is real or complex. In this work, we explore what happens when $\kappa$ is time dependent. According to the adiabatic theorem, the expression of $U_{mn}\left(t\right)$ still holds true for slowly varying $\kappa(t)$ during a short period of time. Considering the case with $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number and the frequency of $\kappa$ is small, $\mathcal{\omega}\ll\mathcal{\omega}_{0}$. It can be predicted that $\left|\Phi(t)\right\rangle$ and $P(t)$ are still approximately periodic even for the quasiadiabatic process. However, this approach could be invalid for the diabatic process. To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes with different values of $\mathcal{\omega}/\mathcal{\omega}_{0}$. We compute the temporal evolution of a site state, $\left|\Phi(0)\right\rangle=|0\rangle$ by using a uniform mesh in the time discretization for the time-dependent Hamiltonian $H(t)$. The evolved state is
@@
where $\mathcal{T}$ is the time-order operator. Fig. 2 shows the plots of $P_{n}(t)$ and $P(t)$ for several typical values of $\mathcal{\omega}/\mathcal{\omega}_{0}$. The numerical results indicate that the Bloch oscillation exists for small $\mathcal{\omega}/\mathcal{\omega}_{0}$. For non-Hermitian cases, the profile of the evolved state does not change substantially, but the probability is periodic or quasiperiodic.
IV Floquet-resonance dynamics
Now, we investigate the dynamics of the system with time-dependent $\kappa(t)$. When $\kappa(t)$ is periodic, the system becomes a Floquet system. The solution of the Schrödinger equation
@@ IV Floquet-resonance dynamics
$\frac{\partial a_{n}}{\partial t}=\frac{1}{\omega_{0}}\frac{\partial\kappa}{\partial t}(a_{n-1}e^{i\mathcal{\omega}_{0}t}-a_{n+1}e^{-i\mathcal{\omega}_{0}t}),$ (21)
which is the starting point for the following discussions. In this work, we consider the case with $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number. Taking the rotating-wave approximation (RWA) under the condition $\left|\mathcal{\omega}-\mathcal{\omega}_{0}\right|\ll\mathcal{\omega}_{0}$, we have
@@
V Dynamical exponents

### v3

@@ Abstract
We study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation, where the hopping strength is a periodic function of time with Floquet frequency $\mathcal{\omega}$. It is shown that the energy level of the instantaneous Hamiltonian is still equally spaced and independent of time $t$ and the Hermiticity of the hopping term. In the case of off resonance, the dynamics are still periodic, while the occupied energy levels spread out at the resonance, exhibiting $t^{z}$ behavior. Analytical analysis and numerical simulation show that the level-spreading dynamics for real and complex hopping strengths exhibit distinct behaviors and are well described by the dynamical exponents $z=1$ and $z=1/2$, respectively.
Keywords: Floquet Hamiltonian, non-Hermitian, resonance, anomalous diffusion, dynamical exponents, 2D simulator
1 Introduction
The Hermitian and static system is the main study object of traditional quantum mechanics, preserving the reality and conservation of energy. When these two constraints are violated, the properties of the material could be drastically modified, leading to new phases and phenomena [1, 2] without analogs in their static or Hermitian counterparts. Floquet engineering has facilitated the exploration of a diverse range of dynamic, topological, and transport phenomena across various physical contexts, including platforms involving ultracold atoms [3, 4, 5, 6]. More recently, the Floquet approach has been employed to engineer topological phases in non-Hermitian systems. Importantly, it was discovered that the interplay between time-periodic driving fields and the presence of gain, loss, or nonreciprocal effects can lead to the emergence of topological phases exclusive to non-Hermitian Floquet systems [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] .
Motivated by the above investigations, the aim of the present paper is to examine how non-Hermiticity and periodic driving impact a Stark ladder system. Specifically, we study the dynamics of the non-Hermitian Floquet Wannier-Stark system in the framework of the tight-binding approximation. A conventional Wannier-Stark ladder consists of quantized equidistant energy levels, which separated by the Bloch frequency and emerge in a periodic Hermitian system under a linear potential [17, 18, 19]. Based on this framework, the Bloch oscillations in cold atoms within the optical lattice have also been theoretically investigated [20, 21, 22].
Recent works [23, 24, 25, 26] show that non-Hermitian hopping terms, including asymmtric and complex strengths, remain the reality of the Wannier-Stark ladder. On the other hand, it has proposed that a time-dependent hopping strength for a Bloch system can be realized by dynamically modulated waveguide arrays [27].
[…]
This paper is organized as follows. In Sec. 2, we introduce a non-Hermitian time-dependent tight-binding model supporting a Wannier-Stark ladder and present the analytical eigenstates of the instantaneous Hamiltonian. In Sec. 3, we focus on the dynamics of the static system, while the Floquet system at resonance in Sec. 4. Sec. 5 describes the application of the solutions to a specific initial state. In Sec. 6, a scheme for the simulator of our results is proposed in the framework of wavepacket dynamics in a $2$D square lattice. Finally, we summarize our results in Sec. 7.
@@ 1 Introduction
Figure 1: (a) Schematic illustrations of the Hamiltonian in Eq. (1), which represents a tight-binding chain with a time dependence on the hopping strength between two adjacent sites. The on-site potential is tilted with slope $\omega_{0}$. It is shown that such a system supports a Stark-ladder with time-independent energy-level spacing. (b) Schematic illustrations of the Hamiltonian in Eq. (48), which is a tight-binding model of a square lattice with time-independent system parameters. The contributions of hopping strength and on-site potential can be used to simulate the dynamics of system (a).
2 Model and generalized Wannier-Stark ladder
@@
where $|n\rangle$ denotes a site state describing the Wannier state localized on the $n$th period of the potential. Here, $\kappa(t)$ is the time-dependent hopping strength, while $\mathcal{\omega}_{0}$ is the slope of the linear potential from a static force. Fig. 1(a) shows a schematic of the system. In our study, $\kappa(t)$ can be real and complex. In order to solve the time-dependent Hamiltonian, we start with the solution for the case where $\kappa$ is an arbitrary complex constant. We then discuss the time-dependent hopping strength $\kappa(t)$ at the end of the Section 3. The non-Hermitian Hamiltonians with complex hoppings have been realized in the coupled waveguides [31], optical fibres [32] and electrical circuit resonators [33].
It is well known that when the hopping constant $\kappa$ is real, the Hamiltonian ${H}$ in momentum space, based on the Bloch basis representation, can be exactly diagonalized. Consequently, the energy levels of ${H}$ are equidistant, which supports the periodic dynamics [34, 35, 36]. This phenomenon has been observed in systems of semiconductor superlattices and ultracold atoms [37, 38, 39, 40]. Now, we extend the conclusion to the case with arbitrary $\kappa$ through an alternative approach. The eigenstate of ${\left|\psi_{m}\right\rangle}$ can always be written in the form
${\left|\psi_{m}\right\rangle=\sum_{n}c_{nm}|n\rangle,}$ (2)
satisfying the Schrödinger equation
${H\left|\psi_{m}\right\rangle=E_{m}\left|\psi_{m}\right\rangle.}$ (3)
The eigenenergy $E_{m}$ can be obtained by the equation
$E_{m}={\kappa}\frac{c_{n+1,m}+c_{n-1,m}}{c_{nm}}+n\omega_{0},$ (4)
which accords with the recurrence relation of the Bessel function $J_{n}$ of argument $-2\kappa/\omega_{0}$, i.e.,
$m\omega_{0}={\kappa}\frac{J_{n-m+1}(-\frac{2\kappa}{\omega_{0}})+J_{n-m-1}(-\frac{2\kappa}{\omega_{0}})}{J_{n-m}(-\frac{2\kappa}{\omega_{0}})}+n\omega_{0}.$ (5)
@@
$E_{m}=m\omega_{0},\ {c_{nm}=J_{n-m}(-\frac{2\kappa}{\omega_{0}}).}$ (6)
We can see that the eigenenergy ${m\mathcal{\omega}_{0}}$ is independent of $\kappa$, regardless of whether it is real or complex. This is crucial for the main conclusion of this work. Although the profile of the eigenstate
@@
${\ \text{IPR}_{m}=\frac{\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{4}}{(\sum\limits_{n}\left|\langle n|\psi_{m}\rangle\right|^{2})^{2}}=\frac{\sum_{n}|J_{n-m}(-\frac{2\kappa}{\omega_{0}})|^{4}}{(\sum_{n}|J_{n-m}(-\frac{2\kappa}{\omega_{0}})|^{2})^{2}}.}$ (8)
@@
Figure 2: Plots of the real and imaginary parts of energy levels for the system at different sizes. The parameters of the system are $\kappa=1+1i$ and $\omega_{0}=1$. One can observe that the real, equally spaced energy levels emerge from the center of the spectrum as the system size increases.
In parallel, without loss of generality, we have
@@
The obtained solution to the non-Hermitian Hamiltonian obeys the well-established theory of non-Hermitian quantum mechanics [41, 42, 43, 44, 45]. In the context of non-Hermitian systems with $H\neq H^{{\dagger}}$, biorthonormal complete eigenstates refer to a set of eigenstates that are biorthonormal and form a complete basis for the Hilbert space of the system. The eigenstates of a non-Hermitian Hamiltonian can be complex and may not be orthogonal to each other. However, if a non-Hermitian Hamiltonian has a complete set of biorthonormal eigenstates, it means that: There exists a left eigenstate $\langle\varphi_{n}|$ and a right eigenstate $|\psi_{n}\rangle$, satisfying
$H|\psi_{n}\rangle=E_{n}|\psi_{n}\rangle,H^{{\dagger}}|\varphi_{n}\rangle=E_{n}^{\ast}|\varphi_{n}\rangle.$ (12)
for each eigenvalue $E_{n}$ such that
$\langle\varphi_{n}|\psi_{m}\rangle=\delta_{nm},\sum_{n}|\psi_{n}\rangle\langle\varphi_{n}|=1.$ (13)
The set of right eigenstates $\left\{|\psi_{n}\rangle\right\}$ and the set of left eigenstates $\left\{|\varphi_{n}\rangle\right\}$ together form a basis for the Hilbert space, allowing any state in the space to be expanded in terms of these eigenstates.
The concept of biorthonormal complete eigenstates is particularly important in non-Hermitian quantum mechanics because it plays an analogous role, both conceptually and computationally, to those in Hermitian quantum systems. For example, when we calculate the time evolution of a state, we first express the initial state $|\psi(0)\rangle$ as a linear combination of the right eigenstates
$|\psi(0)\rangle=\sum_{n}|\psi_{n}\rangle\langle\varphi_{n}|\psi(0)\rangle$ (14)
by projecting the initial state onto the left eigenstates. The state of the system at any later time $t$ is given by
$|\psi(t)\rangle=\sum_{n}|\psi_{n}\rangle e^{-iE_{n}t}\langle\varphi_{n}|\psi(0)\rangle=U(t)|\psi(0)\rangle$ (15)
where $U(t)$ is time evolution operator and will be obtained in the following for the present Hamiltonian in Eq. (1).
Figure 3: Plots of $P_{n}(t)/P(t)$ (see the main text) and $P(t)$ defined in Eqs. (19 ) and (20), obtained by numerical diagonalization for several typical values of system parameters with $\mathcal{\omega}_{0}=1$. Here, (a1), (b1), and (c1) correspond to the cases with $\omega=0$, $0.01$, and $0.1$, respectively, and $\kappa_{0}=1$. (a2), (b2), and (c2) correspond to the cases with $\omega=0$, $0.01$ , and $0.1$, respectively, and $\kappa_{0}=i$. (a3), (b3),and (c3) correspond to the cases with $\omega=0$ , $0.01$, and $0.1$, respectively, and $\kappa_{0}=e^{i\pi/4}$. We can see that the patterns in (a2, b2, c2) and (a3, b3, c3) are similar. This inidcates that the imaginary part of hopping plays an important role in the shapes of evolved states. According to the analytical analysis, ${P(t)}$ is periodic with the period depending on $\omega$ and $\omega_{0}$ When $\omega$ approaches $\omega_{0}$ , the period becomes infinite. The non-periodic behavior observed in (c1, c2, c3) arises because the scale plotted in the figure is smaller than the period.
Figure 4: Plots of $P_{n}(t)/P(t)$ (see the main text) and $\text{In}P(t)$ defined in Eqs. (19 ) and (20), obtained by numerical diagonalization for several typical values of system parameters with $\mathcal{\omega}_{0}=1$. Here, (a1), (b1), and (c1) correspond to the cases with $\omega=0.6$, $1$, and $1.5$, respectively, and $\kappa_{0}=1$. (a2), (b2), and (c2) correspond to the cases with $\omega=0.6$, $1$ , and $1.5$, respectively, and $\kappa_{0}=i$. (a3), (b3),and (c3) correspond to the cases with $\omega=0.6$, $1$, and $1.5$, respectively, and $\kappa_{0}=e^{i\pi/4}$.
3 Bloch oscillation in a non-Hermitian system
The dynamics of the system are governed by the propagator
@@ 3 Bloch oscillation in a non-Hermitian system
the probability distribution in real space for the evolved state $\left|\Phi(t)\right\rangle$ is
@@
which is periodic. We obtain the conclusion that the total probability
@@
is still periodic, regardless of whether $\kappa$ is real or complex. In practice, the measurement of $P_{n}(t)$ has been reported in many non-Hermitian experiments [46, 47, 48, 49].
Up until now, we have been considering $\kappa(t)=\kappa$ as time-independent; now, we explore what happens when $\kappa(t)$ is time dependent. According to the adiabatic theorem, the expression of $U_{mn}\left(t\right)$ still holds true for slowly varying $\kappa(t)$ during a short period of time. Considering the case with $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number and the frequency of $\kappa$ is small, $\mathcal{\omega}\ll\mathcal{\omega}_{0}$. It can be predicted that $\left|\Phi(t)\right\rangle$ and $P(t)$ are still approximately periodic even for the quasiadiabatic process. However, this approach could be invalid for the diabatic process.
To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes with different values of $\mathcal{\omega}/\mathcal{\omega}_{0}$. We compute the temporal evolution of a site state, $\left|\Phi(0)\right\rangle=|0\rangle$ by using a uniform mesh in the time discretization for the time-dependent Hamiltonian $H(t)$. The evolved state is
@@
where $\mathcal{T}$ is the time-order operator. Figs. 3 and 4 are the plots of $P_{n}(t)/P(t)$ and $P(t)$ for several typical values of $\omega/\omega_{0}$. For the non-Hermitian systems, the difference between the maximum and minimum of $P(t)$ is very large. Thus the profile of the evolved state cannot be properly presented by the plot of $P_{n}(t)$. To present a complete and visible profile, we plot the rescaled $P_{n}$ by $P_{n}(t)/P(t)$. When the total probability $P(t)$ is periodic in time, the periodicity of the evolved state can be confirmed by the profile of the probability distribution.
The numerical results indicate that the Bloch oscillation exists for small $\mathcal{\omega}/\mathcal{\omega}_{0}$. For non-Hermitian cases, the profile of the evolved state does not change substantially, but the probability is periodic. From Fig. 4 we can see that the evolution of the profile of the evolved state is still periodic when $\omega\mathbf{\neq}\omega_{0}$, but diffuses outward at $\omega=\omega_{0}$. This indicates that the dynamics are peculiar at resonance, which is the focus of the paper.
4 Floquet-resonance dynamics
Now, we investigate the dynamics of the system with time-dependent $\kappa(t)$. As previously mentioned, we consider periodically modulated $\kappa(t)=\kappa_{0}\cos(\omega t)$, where $\kappa_{0}$ can be a complex number. The solution of the Schrödinger equation
@@ 4 Floquet-resonance dynamics
$\frac{\partial a_{n}}{\partial t}$ $=$ $\frac{1}{\omega_{0}}\frac{\partial\kappa}{\partial t}(a_{n-1}e^{i\mathcal{\omega}_{0}t}-a_{n+1}e^{-i\mathcal{\omega}_{0}t})$ (25)
$=$ $-\frac{{\kappa_{0}}\omega}{\omega_{0}}{\sin(\omega t)}(a_{n-1}e^{i\mathcal{\omega}_{0}t}-a_{n+1}e^{-i\mathcal{\omega}_{0}t}),$
which is the starting point for the following discussions.
Taking the rotating-wave approximation (RWA) under the condition $\left|\mathcal{\omega}-\mathcal{\omega}_{0}\right|\ll\mathcal{\omega}_{0}$, we have
@@
lt clearly shows that the eigenvalues of the equivalent Hamiltonian are proportional to the hopping constant $\kappa_{0}$, and then determines the behavior of the amplitude of the evolved state which can be conservative, damping, or explosive, respectively. The real and imaginary parts of $\kappa_{0}$ will affect the real and imaginary nature of the energy levels, which in turn will affect whether the probability is conserved and the evolution of the wave function. It can be anticipated that the presence of an imaginary part in $\kappa_{0}$ will significantly influence the dynamical evolution of the system.
5 Dynamical exponents